Prove: cot30° sin30° cos30° = 1-sinē30°
Answer by jim_thompson5910(35256)

cot(30)*sin(30)*cos(30) = 1-sin^2(30)

(cos(30)/sin(30))*sin(30)*cos(30) = 1-sin^2(30)

cos(30)*cos(30) = 1-sin^2(30)

cos^2(30) = 1-sin^2(30)

1-sin^2(30) = 1-sin^2(30)

So the identity has been verified.